# SensorChat: Answering Qualitative and Quantitative Questions during Long-Term Multimodal Sensor Interactions

### v1

@@ 3.1. Problem Statement
We target real-world QA scenarios in which users carry mobile sensing devices during their daily routines, with minimal attention to the sensing process. In this context, sensor data is collected continuously over an extended period in natural, uncontrolled environments. During this process, users may pose arbitrary, personalized questions about the sensor data, ranging from inquiries about a single day to those spanning several weeks, and expect precise answers derived from the collected data.
Formally, the problem is formulated as generating a proper answer given the question and sensory inputs, denoted as $\mathbf{x}$. The questions can be arbitrary, including qualitative and quantitative queries, while the answers are expected to accurately address users’ concerns. The key distinction between our problem formulation and existing QA approaches lies in the format of $\mathbf{x}$, which consists of long-duration high-dimensional data collected up to the moment the question is asked. As shown in Table 1, this makes our problem significantly more challenging.
3.2. Limitations of SOTA Methods
@@
With a concrete problem setup, we conduct a motivating study to understand the limitations of state-of-the-art systems on this problem. We then discuss the lessons learned which inspire the design of SensorChat.
Dataset Setup We choose the SensorQA dataset [^1] https://github.com/benjamin-reichman/SensorQA (Reichman et al., 2025) to simulate long-term real-world QA interactions between humans and sensing devices. To our knowledge, SensorQA is the first and only existing QA dataset that matches our target scenario, i.e., in-the-wild, long-term, time series sensor data collection with practical questions that align with users’ interests.
Specifically, SensorQA (Reichman et al., 2025) used the ExtraSensory dataset (Vaizman et al., 2017, 2018b) as its sensor data source, which contains raw sensor measurements from 60 subjects and has more than 300K minutes of data, annotated with 51 activity or context labels after cleaning. The data were collected using easily accessible smartphone and smartwatch sensors, including IMUs, audio (MFCC features), and phone state sensors (compass, location, Wi-Fi, light, etc.). Unlike other human activity recognition datasets (Reyes-Ortiz et al., 2012; Banos et al., 2014; Roggen et al., 2012; Ouyang et al., 2023), ExtraSensory is notable for its natural and uncontrolled collection environment, with no restrictions on device placement.
The QA data in SensorQA (Reichman et al., 2025) were generated by real human workers as if they were interacting with a smart sensor device about their daily lives. As a result, SensorQA (Reichman et al., 2025) comprises diverse and practical questions that provide valuable insights for users. SensorQA covers a wide range of topics from normal activities to work patterns to work-life balance, and question types such as time queries, counting, and activity recognition. In SensorQA (Reichman et al., 2025), the authors identified six question categories and seven answer categories. Table 2 details the categories along with example Q&As and the corresponding number of instances. While SensorQA exhibits a skewed distribution favoring quantitative and time-related queries, this trend underscores users’ strong interest in gaining temporal insights into their lives, highlighting the need for a QA system to address quantitative queries.
Question Example Questions # of
Categories Questions
Time Compare Did I spend more time sitting or standing? 1,432
Day Query On which day did I spend the most time at home? 1,277
Time Query How long was I in class and at school? 1,119
Counting How often did I groom? 725
Existence Did I have a meeting on Wednesday? 668
Action Query What did I do after I left home on Tuesday? 428
@@
Answer Example Shortened # of
Categories Answers Answers
Action Doing computer work 1,357
Day/Days Last Friday 1,242
Existence Yes/No 1,047
Time Length 40 Minutes 1,018
Location At school 792
Count Three times 401
Timestamp Around 11:00 am 310
@@
Table 2. Q&A categories in the SensorQA dataset (Reichman et al., 2025). The short answers are presented for simplicity.
In total, SensorQA (Reichman et al., 2025) contains 5.6K QA pairs reflecting practical human interests in daily life monitoring, including about 3K questions on daily activities and 2.6K on longer durations spanning weeks or months. We randomly split the QA data in SensorQA (Reichman et al., 2025) into an 80/20 train-test set. During the experiments, the full sensor data history for each user and one question are fed to the baselines. The generated answer is compared with the ground-truth answer in SensorQA (Reichman et al., 2025).
State-of-the-art (SOTA) Baselines Setup In this motivating study, we select three state-of-the-art methods that have open-sourced code, each representing one mainstream category of approaches:
  • DeepSQA (Xing et al., 2021) represents the SOTA CNN-LSTM based model for fusing the question and time series sensor data. We specifically train and test the DeepSQA-CA model which uses Compositional Attention with CNN-LSTM architecture and achieves the highest accuracy as reported in the original paper. We use full-history IMU signals as sensor inputs, since the DeepSQA-CA model was designed to handle IMU time series data.
  • IMU2CLIP+GPT-4 (Moon et al., 2023b) is the SOTA GPT-based generative method. It uses IMU2CLIP (Moon et al., 2023b) to associate raw IMU signals with text and uses GPT-4 for reasoning and answer generation. For each timestamp, IMU2CLIP searches for a text description that best matches the IMU data. The complete narrative text from all timestamps is then fed into GPT-4 (Achiam et al., 2023) with the question to generate an answer. We train the IMU2CLIP model using the IMU sensor data in SensorQA (Reichman et al., 2025).
  • LLaMA-Adapter (Zhang et al., 2023) is the SOTA multimodal LLM model that integrates image or video inputs with the LLaMA backbone using a Transformer-based adapter. We use the activity Gantt charts as image inputs, representing the sensor information. We start with their pretrained weights on a LLaMA2-7B (Touvron et al., 2023a) backbone and then finetune it with LoRA (Hu et al., 2021) on SensorQA (Reichman et al., 2025).
(a) Example of single-day query.
(b) Example of multi-day query.
Figure 2. Ground-truth QA pairs in SensorQA (Reichman et al., 2025) and SOTA methods’ answers.
Lesson 1: LLMs are necessary for natural interactions, but not enough We first visualize two examples of the question, true answer and the response generated by the three SOTA methods in Fig. 2 (a) and (b). Similar as in Fig. 1, we visualize the user’s activity in Gantt charts. In Fig. 2 (a), we study a single-day query asking the activities after waking up on a specific day of Wednesday. While the answers are obvious to humans from the chart, this question requires multi-step reasoning: the algorithm first needs to figure out the end time of ”waking up”, and then queries the activity right after this timestamp. As a result, traditional methods such as DeepSQA cannot follow these non-trivial reasoning steps needed to process the full history of time series data. Additionally, DeepSQA can only generate short phrases because the CNN-LSTM is trained for classification, learning to select the most suitable answer from a limited set of candidates. The other two methods based on LLMs are able to produce complete sentences, showing LLM’s crucial role in natural interactions with humans.
However, LLMs do not always succeed. In Fig. 2 (a), LLaMA-Adapter ends up with an irrelevant answer of ”went to school”, likely a random guess rather than a result of multimodal fusion. Multimodal LLMs like LLaMA-Adapter may underperform without extensive high-quality data (e.g., over 1M samples) for finetuning. In this example, only IMU2CLIP+GPT-4 proves its ability to generate the correct answer of ”showering” and ”in the toilet”. This can be attributed to IMU2CLIP’s success in associating the IMU signals with the correct text, and GPT-4 generating the correct reasoning from the full narrative text. Nevertheless, IMU2CLIP+GPT-4 has a significant limitation due to GPT’s restricted prompt length. This leads to limited narrative text and a reduced sensor duration that can be processed. Specifically, IMU2CLIP+GPT-4 struggles with multi-day queries, as shown in Fig. 2 (b). The question asks for comparing the frequency of meals on multiple days, and then identifying the day with the fewest meals. All three methods fail in this challenging task. IMU2CLIP+GPT-4 in particular refuses to give any valid answer.
Lesson 2: SOTA methods struggle with accurate quantitative answers over long-duration sensor data Although we recognize the critical role of LLMs in user interactions, it is well known that LLMs struggle with quantitative analyzes and numerical reasoning, especially for long-context inputs (Li et al., 2024; Gu and Dao, 2023). We study the precision of answers generated by SOTA methods on SensorQA (Reichman et al., 2025). To measure accuracy, we match the key phrases from the ground-truth answers with those in the generated answers. Specifically, SensorQA offers a shortened version of each answer by distilling 1-2 key words from the full answers using GPT-3.5-Turbo (Ye et al., 2023). A generated answer is considered correct if it includes the ground-truth short answer; otherwise, it is marked as incorrect.
@@
Figure 3. Answer accuracy of state-of-the-art methods and SensorChat by question category (left) and answer category (right). Categories are arranged from easier questions/answers on the left to more challenging questions/answers on the right.
Fig. 3 shows the accuracy results of SOTA methods by question category (left) and answer category (right). One key observation is that different Q&A types present varying levels of difficulty for the baselines. For example, existence and time comparison questions are easier to answer due to the limited number of possible responses (e.g., Yes/No for existence questions). Counting and time queries are more challenging as they require quantitative analysis. Additionally, different methods exhibit varying sensitivity to question and answer types. IMU2CLIP+GPT-4 completely fails in counting questions, as highlighted by the red arrow in Fig. 3 (left). This can be credited to GPT-4’s difficulty with counting in long narrative texts. Similarly, as pointed by the red arrow in the right figure of Fig. 3, all three baselines struggle to generate correct answers for timestamps, underscoring the weaknesses of existing methods in quantitative reasoning over long-duration sensor data. This weakness stems from the essence of existing designs, where long-duration raw sensor signals are directly fed into large models for fusion with text. Our key insight is that sensor-text fusion requires a more carefully designed approach to handle the diverse queries and long-duration multimodal sensor data encountered in real-life scenarios. To address this, we propose SensorChat, a novel system with a three-stage pipeline that includes a dedicated sensor data query stage. As shown in Fig. 3, SensorChat significantly improves answer accuracy across all categories, particularly for quantitative tasks such as day comparisons and time queries.
@@ 4.1. Overview of SensorChat
To address the challenge of accurate QA over long-duration sensor data, we design a novel three-stage pipeline in SensorChat, consisting of question decomposition, sensor data query, and answer assembly. Specifically, we use LLMs in question decomposition and answer assembly, recognizing their essential roles in correctly interpreting user queries and generating natural language answers. The intermediate sensor query stage is the key component of SensorChat. First, the query stage uses a pretrained sensor encoder to effectively encode high-dimensional and long-duration multimodal sensor timeseries into meaningful embeddings. Second, the query stage performs a similarity search in the embedding space to retrieve all sensor information relevant to the original question. This ensures precise extraction of sensor context, making a significant contribution to the accuracy of the final answer. To the best of the authors’ knowledge, SensorChat is the first system to incorporate an explicit sensor data query stage for accurately handling sensor-based language tasks in long-term monitoring.
@@
The complete pipeline of SensorChat is illustrated in Fig. 4. For example, given a question like, “How long did I exercise last week in the morning?” (① in Fig. 4), SensorChat first decomposes the question and generates specific sensor data queries using LLMs. These queries include details such as the context of interest, date and time ranges, and the summarization function (③). In this case, the query might specify a context of “do exercise,” a time span of “last week,” a time of day of “morning,” and a summarization function of CalculateDuration. To enhance the precision of question decomposition, SensorChat uses solution templates (②) with carefully designed prompts. Next, the sensor query stage encodes the activity text “exercise” into the embedding space (④) and retrieves sensor embeddings that are sufficiently similar to the text embedding (⑤). These sensor embeddings are encoded offline from the full-history raw sensor data. Both the sensor and text encoders are pretrained offline to align their outputs in the same embedding space, ensuring accurate sensor data retrieval. Therefore, only text encoding and similarity-based query searches are performed online. Additional properties, such as the date (“last week”) and time of day (“morning”), are used to constrain the sensor query range. The final step of the sensor data query stage involves summarizing the relevant sensor embeddings into a textual context with a summarization function (⑥). For instance, with the CalculateDuration function, the summarization step calculates the total duration of the retrieved sensor embeddings. This results in a sensor context such as: “Among all days last week, you exercised for 35 minutes on Monday morning and 55 minutes on Thursday morning.” The summarization function is identified during question decomposition. Finally, the original question and the sensor context are fed into the answer assembly stage, where a fine-tuned LLM generates the final answer (⑦). With the precise sensor context, the model produces an accurate response, such as: “You exercised a total of 1 hour and 30 minutes last week.” (⑧).
Figure 4. The system diagram of SensorChat including three stages.
In the following lines, we describe the key design elements of SensorChat: the offline contrastive sensor-text encoder pretraining (Sec. 4.2) and the online three-stage pipeline (Sec. 4.3).
4.2. Contrastive Sensor-Text Pretraining for Partial Contexts
The sensor and text encoders are crucial for effective sensor-text fusion, serving as a ”bridge” between high-dimensional sensor timeseries and semantic text. A high-quality sensor encoder ensures that SensorChat can accurately and comprehensively retrieve relevant sensor embeddings during the sensor data query stage. Similarly, a robust text encoder is essential for handling the arbitrary contextual text generated during the question decomposition stage.
However, achieving this level of alignment poses significant challenges due to the lack of suitable techniques. Existing pretraining methods, such as CLIP and its variants (Radford et al., 2021; Moon et al., 2023b), are effective for pretraining encoders with paired inputs, i.e., one piece of sensor data and one sentence. Unfortunately, encoders trained in this manner struggle to accurately identify similar sensor embeddings when provided with partial or arbitrary context instead of complete sentences. For instance, using CLIP, a piece of sensor data might align well with the full sentence ”The person is sitting and working on computers at school.” However, when a user is specifically interested in a partial context like ”working on computers,” the encoded text embedding may not closely match the original sensor embeddings, leading to reduced accuracy in sensor data query as we show in Sec. 6.6. To address this challenge, we introduce a novel contrastive sensor-text pretraining loss for partial contexts.

### v2

@@ 3.1. Problem Statement
We formulate the problem as generating an appropriate answer based on a given question and sensor input. In the $i$th interaction round, the input and output are a pair of natural language question and answer, denoted as $(q^{i},a^{i})$ for $i\in{1,...,N}$, where $N$ is the total number of QA interactions. The sensor input consists of the full history of raw sensor signals $\mathbf{x}\in\mathbb{R}^{T\times d}$, where $T$ is the number of recorded time steps and $d$ is the number of multimodal sensors. We specifically target long-duration, high-frequency sensor data (e.g., raw IMU signals sampled at 40Hz over more than a day), resulting in extremely large $T$. Formally, our objective is formed as training an autoregressive model, $\omega$, to predict the next token in the answer sequence:
(1) $\max_{\omega}\sum_{i=1}^{N}\sum_{j=1}^{m}\log p_{\omega}\left(a^{i}_{j}\>{|}\>\mathbf{x},q^{i},a^{i}_{1:j-1}\right),$
where $i$ iterates through all QA pairs, $j$ iterates through all tokens in the answer and $m$ is the maximal length for answers. Given a QA dataset $(q^{i},a^{i})$, $\omega$ is trained to maximize the likelihood of the correct next token. During inference, $\omega$ is fixed, and the model selects the next token with the highest predicted probability.
Question and Answer Types We target at two broad categories of user queries $q^{i}$: quantitative and qualitative questions, as commonly identified in prior work (Xing et al., 2021; Reichman et al., 2025; Henrich and Herschbach, 2000). For quantitative questions, we mainly focus on the six question types and seven answer types defined by the SensorQA dataset (Reichman et al., 2025), a state-of-the-art quantitative QA benchmark for daily-life monitoring. Table 3 shows the detailed question and answer types. SensorQA is chosen for its realism, as its questions were written by human annotators simulating interactions with a smart sensor system. Due to the lack of a benchmark for qualitative QA using long-term sensor data, we focus on general topics that affect quality of life, such as work–life balance and social interactions (Henrich and Herschbach, 2000). We evaluate SensorChat on qualitative questions through a real-user study in Sec. 7.
3.2. Limitations of State-of-the-Art (SOTA) Methods
@@
With a concrete problem setup, we conduct a motivating study to understand the limitations of state-of-the-art systems on this problem. We then discuss the takeaways which inspire the design of SensorChat.
Study Setup We select the SensorQA dataset [^1] https://github.com/benjamin-reichman/SensorQA (Reichman et al., 2025) because it has the best alignment with QA scenarios using long-duration, high-frequency multimodal sensor data. SensorQA has raw sensor measurements collected using smartphone and smartwatch sensors, including IMUs, audio (MFCC features), and phone state sensors (compass, GPS location, Wi-Fi, light, etc.). It contains data from 60 subjects and has more than 300K minutes of data, annotated with 51 context labels after cleaning. In total, SensorQA (Reichman et al., 2025) contains 5.6K quantitative QA pairs reflecting practical human interests in daily life monitoring, including approximately 3K questions on a single day and 2.6K for longer durations spanning weeks or months. We randomly split the data in SensorQA (Reichman et al., 2025) into an 80/20 train-test set. Raw sensor data is split into time windows of 20 seconds for models to process.
Question Categories Example Questions
Time Compare Did I spend more time sitting or standing?
Day Query On which day did I spend the most time at home?
Time Query How long was I in class and at school?
Counting How often did I groom?
Existence Did I have a meeting on Wednesday?
Action Query What did I do after I left home on Tuesday?
@@
Answer Categogies Example Answers
Action Doing computer work
Day/Days Last Friday
Existence Yes/No
Time Length 40 Minutes
Location At school
Count Three times
Timestamp Around 11:00 am
@@
Table 3. Quantitative Q&A categories that we focus on in this paper, as identified in the SensorQA dataset (Reichman et al., 2025). The shorted version of the answers are presented for simplicity.
We train and evaluate three methods representing SOTA solutions using neural network and LLM-based models: (1) LLaMA2-7B (Touvron et al., 2023b), as a reference, is the LLM baseline that only considers the question without sensor data. (2) DeepSQA (Xing et al., 2021) represents the SOTA CNN-LSTM based model for fusing the question and time series sensor data. While DeepSQA was originally designed for short-duration sensor data, we adapt it by using the full history of IMU signals as sensor-side input. (3) LLMSense (Ouyang and Srivastava, 2024) is the SOTA LLM-based method for reasoning over sensor data. It converts neural network classifications into a narrative log including all timestamps, and then feeds this log and the user question into an LLM to generate an answer. Detailed implementations of these methods are discussed in Appendix D. We focus on answer accuracy for quantitative questions in SensorQA. An answer is considered correct if it contains the key phrases from the ground-truth answer provided by SensorQA. Otherwise, it is marked as incorrect. For example, in Fig. 1(c), if the ground-truth answer is “You spent the longest eating time on Saturday,” then an answer that includes ”Saturday”, such as SensorChat’s answer, is correct. An answer like ”Monday” is incorrect.
Refer to caption Refer to caption
(a) Question categories (b) Answer categories
@@
Figure 2. Answer accuracy of state-of-the-art methods and SensorChat on SensorQA (Reichman et al., 2025). Categories are arranged from easier questions/answers on the left to more challenging questions/answers on the right. Red arrows highlight the question/answer types that existing works struggle with.
Fig. 2 shows the accuracy results of SOTA methods by question category (left) and answer category (right). The key observation is that all SOTA baselines deliver low answer accuracy. LLaMA2-7B (Touvron et al., 2023b), DeepSQA (Xing et al., 2021), and LLMSense (Ouyang and Srivastava, 2024) achieve average accuracies of 0.26, 0.27, and 0.23, respectively. Notably, DeepSQA and LLMSense—both of which attempt to integrate sensor data with the question—perform comparably to or worse than LLaMA2-7B, which does not incorporate any sensor input. This suggests that existing approaches struggle to handle the complexity introduced by long-duration, high-frequency sensor data. In particular, DeepSQA’s performance suffers when directly fed with more than a day’s worth of IMU signals. The increased input size and complexity significantly complicate the fusion process, overwhelming its relatively simple model architecture. On the other end, LLMSense performs poorly on questions requiring precise numerical reasoning, such as counting or time queries (highlighted by red arrows in Figure 2). This is due to LLMSense’s reliance on LLMs to reason over long narrative logs, which contains up to 1,440 lines for a single day. Prior studies have shown that LLMs often struggle with long-context numerical tasks and are prone to hallucinations in such settings (Huang et al., 2023; Gu and Dao, 2023; Li et al., 2024).
In summary, directly applying existing QA methods to long-duration, high-frequency sensor data leads to poor performance, particularly in quantitative reasoning tasks. These limitations highlight the need for a new system capable of effectively processing raw sensor inputs at scale and fusing them with practical user queries. To address this, we propose SensorChat, a novel system with a three-stage pipeline that includes a dedicated sensor data query stage. As shown in Fig. 2, SensorChat significantly improves answer accuracy across all categories, particularly for quantitative tasks such as day comparisons and time queries.
@@ 4.1. Overview of SensorChat
We design SensorChat, a novel end-to-end QA system for answering both quantitative and qualitative questions using long-duration, high-frequency raw sensor data. SensorChat achieves the goal by having a three-stage pipeline: question decomposition, sensor data query, and answer assembly. SensorChat addresses all three challenges discussed in Sec. 1, i.e., processing long-duration, high-frequency sensor data, effectively fusing sensor data with text questions, and generating accurate, semantically meaningful answers. The sensor query stage is the key novel component of SensorChat. In the query stage, SensorChat pretrains a sensor encoder and a text encoder offline to effectively associate high-frequency and long-duration multimodal sensor timeseries with semantically meaningful embeddings. To effectively fuse sensor and text modalities, SensorChat further designs a custom online query mechanism using a similarity search, retrieving all sensor information relevant to the original question. This enables precise extraction of sensor context across varying durations and question types, significantly improving the accuracy of the final answer. Last but not least, SensorChat utilizes two LLMs: one in the question decomposition stage to interpret user queries, and the other in the answer assembly stage to generate the final response, further enhanced through in-context learning and fine-tuning techniques.
Figure 3. The system diagram of SensorChat including three stages. When a user asks a question, SensorChat first decomposes the question (stage 1), queries the relevant sensor embeddings (stage 2), and generates accurate, semantically meaningful answers (stage 3). The sensor and text encoder used in the query stage are pretrained offline in advance.
@@
The complete pipeline of SensorChat is illustrated in Fig. 3. For example, given a question like, “How long did I exercise last week in the morning?” (① in Fig. 3), SensorChat first decomposes the question and generates specific sensor data queries using LLMs. These queries include details such as the context of interest, date and time ranges, and the query function (③). In this case, the query might specify a context of “do exercise,” a time span of “last week,” a time of day of “morning,” and a query function of CalculateDuration. To improve the accuracy of question decomposition, SensorChat uses solution templates for in-context learning (②) with carefully designed prompts. In-context learning is chosen because it enhances the quality of LLM outputs using only a small number of solution templates, compared to finetuning (Fu et al., 2023). Next, the sensor query stage encodes the context label “exercise” into a text embedding (④) and retrieves sensor embeddings that are sufficiently similar to the text embedding (⑤). These sensor embeddings are encoded offline from the full-history raw sensor data. Both the sensor and text encoders are pretrained offline to align their outputs in the same embedding space, ensuring accurate sensor data retrieval. Therefore, only text encoding and similarity-based query searches are performed online. Additional properties, such as the date (“last week”) and the time of day (“morning”), are used to constrain the query range. The final step of the sensor data query stage involves summarizing the relevant sensor embeddings into text using the corresponding query function (⑥). For instance, with the CalculateDuration function, this step calculates the total duration of the retrieved sensor embeddings. This results in a sensor context such as: “Among all days last week, you exercised for 35 minutes on Monday morning and 55 minutes on Thursday morning.” Finally, the original question and the sensor context are sent to the answer assembly stage, where a fine-tuned LLM generates the final answer (⑦). Finetuning is used to align the model’s output with human-generated answers from state-of-the-art datasets. With accurate sensor context, the model generates a precise response, such as: “You exercised a total of 1 hour and 30 minutes last week.” (⑧). In the following lines, we describe the detailed designs in each stage of SensorChat: question decomposition (Sec. 4.2), sensor data query (Sec. 4.3) and answer assembly (Sec. 4.4)).
4.2. Question Decomposition
The question decomposition stage processes the user question and identifies specific query triggers for the sensor data query stage, such as the context of interest, date and time ranges, and the query function, as illustrated in the leftmost box in Fig. 3. We choose to rely on LLMs for this stage due to their outstanding capability in handling natural language inputs. However, LLMs are not without limitations as they can produce erroneous outputs or hallucinations (Huang et al., 2023). The key challenge in designing this stage lies in generating accurate decompositions for arbitrary user inputs, particularly in the absence of a dedicated dataset for this task. To address this, SensorChat leverages the capabilities of a pre-trained LLM by employing in-context learning (Shao et al., 2023) and chain-of-thought techniques (Chu et al., 2024). These approaches are chosen because, compared to finetuning (Hu et al., 2021), they enhance the quality of LLM outputs using a small number of manually crafted examples, eliminating the need for a high-quality, task-specific dataset.
An example prompt for question decomposition is illustrated in Fig. 4(a). We instruct the LLM to mark different arguments with distinct symbols, such as, using ”¡¡” and ”¿¿” for function names, ”((” and ”))” for context labels, etc. This enables accurate extraction of multiple arguments from a single LLM interaction. SensorChat also supports multiple extracted phrases to handle complex questions such as “How long did I work at school on Monday and Tuesday?” The question decomposition stage is designed with the flexibility to support new extractable properties. We next explain more details on the in-context learning and chain-of-thought techniques.
In-Context Learning (ICL) integrates a few examples directly into the prompt during inference, enabling LLMs to adapt effectively to specific tasks without requiring fine-tuning (Alayrac et al., 2022; Shao et al., 2023). Building on this insight, we design a library of general solution templates covering diverse QA scenarios and incorporate them as in-context examples to improve decomposition accuracy (Fig. 4(a)). We observe that decomposition solutions for the same question type often share similarities. For instance, ”how long” questions typically map to the CalculateDuration function. To utilize this question-specific property, in SensorChat, we first classify the question type using a BERT model (Devlin et al., 2018). Then, we dynamically select templates that best match the question category. This strategy has been shown to improve ICL performance, even in settings that are different from ours but face the same challenge of limited annotated data (Fu et al., 2023). Fig. 4(b) illustrates an example of a solution template for a time query.
Chain-of-Thought (CoT) Prompting explicitly requests LLMs to generate their reasoning process, which improves accuracy in zero- or few-shot logical reasoning (Chu et al., 2024; Lu et al., 2022). In SensorChat, we include “please generate step-by-step explanations” in prompts to encourage CoT, as shown in Fig. 4(a). CoT and detailed solution templates enhance logical reasoning, especially for multi-step questions.
(a) The prompt design for the question decomposition stage includes ICL and CoT. A list of available query functions, solution templates, and the user’s question need to be included in the prompt.
(b) Example solution template for a time-related query to be inserted into the prompt to question decomposition.
Figure 4. Prompts design in the question decomposition stage of SensorChat to enhance the generating quality of LLMs.
4.3. Sensor Data Query
The sensor query stage, as the most important module in SensorChat, is responsible for processing the long-duration, high-frequency raw sensor signals and extracting the most relevant information given arbitrary user queries. This enables effective fusion of sensor data with natural language. As shown in the middle part of Fig. 3, the sensor data query stage takes the extracted arguments from question decomposition, such as the context of interest (e.g., “do exercise”), time span (e.g., “last week”), time of day (e.g., “morning”) and query function (e.g., CalculateDuration). The output of the stage is a concise summary of the relevant sensor data in text form. SensorChat introduces two novel components to addresses these challenges. First, during offline training, SensorChat pretrains a sensor encoder and a text encoder using a novel contrastive loss to align sensor signals with partial context (Sec. 4.3.1). Then, during online querying, SensorChat designs a set of custom query functions to extract relevant sensor embeddings, which are converted to textual summarizations for the final answer assembly stage (Sec. 4.3.2).
4.3.1. Contrastive Sensor-Text Pretraining for Partial Contexts
The sensor and text encoders are crucial for effective sensor-text fusion, serving as a ”bridge” between high-frequency sensor timeseries and semantic text. A key challenge in SensorChat is ensuring that the sensor and text encoders can handle diverse context phrases extracted from user questions, which may vary unpredictably based on users’ habits. Suppose $c_{i}$ represents the context argument extracted from question $q_{i}$. Since SensorChat is designed for general daily life monitoring, $c_{i}$ can refer to anything from activities, locations, or social interactions (e.g., spending time with family)., depending on the user’s interest. For example, in the question “Did I go to school?”, $c_{i}$ is “go to school,” which refers to a location. This means the encoders in SensorChat must be able to align sensor data with a broad range of contexts. However, achieving this level of alignment is difficult. Existing pretraining methods like CLIP and its variants (Radford et al., 2021; Moon et al., 2023b) work well when aligning sensor data with full, detailed sentences (e.g., “The person is sitting and working on computers at school.”). However, when a user is specifically interested in a partial context like ”working on computers,” the encoded text embedding may not closely match the original sensor embeddings, leading to reduced accuracy in sensor data query as we show in Appendix F. To address this challenge, we introduce a novel contrastive sensor-text pretraining loss for partial contexts.